OpenStudy (anonymous):

xy" -y'+xy' + y = x^2e^(2x) I tried using Euler-Cauchy equation formula, but that doesn't apply to this, or does it ?

OpenStudy (amistre64):

Cauchy does what for us?

OpenStudy (amistre64):

\[xy" -y'+xy' + y = x^2e^{2x} \] \[y" -\frac1xy'+y' +\frac1x y = xe^{2x} \] \[y" +(1-\frac1x)y' +\frac1x y = xe^{2x} \]

OpenStudy (amistre64):

hard to remember much past this :) what would be next? Have you learnted Wronskians yet?

OpenStudy (anonymous):

I have learned Wronskians

OpenStudy (anonymous):

I tried dividing by x as you did, then replacing y= m or r making Y'' = r^2 or m^2

OpenStudy (anonymous):

Basically, how to find y (c) = x and y (p)= x

OpenStudy (amistre64):

\[r^2+(1-\frac1x)r +\frac1x = 0\] \[r=\frac{-(1-\frac1x)\pm\sqrt{(1-\frac1x)^2-\frac4x}}{2}\] just wondering

OpenStudy (amistre64):

im not sure if my idea is proper for "not constant" coeffs

OpenStudy (amistre64):

im thinking of trying a power series solution for the homogenous parts :)

OpenStudy (anonymous):

I will try the power series, but hmmm,,,,

OpenStudy (amistre64):

\[\frac1xy=\sum_0a_nx^{n-1}\]\[(1-\frac1x)y'=\sum_1a_n*nx^{n-1}-\sum_1a_n*nx^n\]\[y''=\sum_2a_n*n(n-1)x^{n-2}\] right?

OpenStudy (amistre64):

adjust things to the smallest exponent, the feed out till we have like indexes

OpenStudy (anonymous):

I'm working on the power series, had to refresh my memory as I've been on variatino of parameters for a while. Thank you ! Keep it coming, if you can.

OpenStudy (amistre64):

\[\sum_{1}a_{n-1}x^{n-2}\\ \sum_2a_{n-1}*(n-1)x^{n-2}-\sum_3 a_{n-2}(n-2)x^{n-2}\\\sum_2a_n*n(n-1)x^{n-2}\] \[a_0~x^{-1}+\sum_{2}a_{n-1}x^{n-2}\] hmm, i wonder if the -1 exponent casues troubles since a poly only has positive exponents

OpenStudy (anonymous):

perhaps we can bring the exponents to positive only,

OpenStudy (anonymous):

times it by

OpenStudy (amistre64):

\[a_0x^{-1}+a_1+\sum_{3}a_{n-1}x^{n-2}\\ a_1+\sum_3 a_{n-1}(n-1)x^{n-2}-\sum_3 a_{n-2}(n-2)x^{n-2}\\2a_2+\sum_3a_nn(n-1)x^{n-2}\] \[a_0x^{-1}+2a_1+2a_2\\~~~+\sum_{3}~[a_{n-1}+a_{n-1}(n-1)-a_{n-2}(n-2)+a_nn(n-1)x^{n-2}]~x^{n-2}\] now this is equal to zero only if all the constants (an) are zero; which makes be think that the x^-1 really doesnt matter. but then again it might :) \[a_{n-1}+a_{n-1}(n-1)-a_{n-2}(n-2)+a_nn(n-1)x^{n-2}=0~:~n=0,1,2,3,...\]

OpenStudy (amistre64):

multiplying everything thru by x doesnt alter the recipe for the ans does it?

OpenStudy (amistre64):

\[a_{n-1}+a_{n-1}(n-1)-a_{n-2}(n-2)+a_nn(n-1)=0\] \[a_n=-\frac{a_{n-1}+a_{n-1}(n-1)-a_{n-2}(n-2)}{n(n-1)}\] \[a_2=-\frac{a_{2-1}+a_{2-1}(2-1)-a_{2-2}(2-2)}{2(2-1)}\]\[a_2=-\frac{2a_{1}}{2}=-a_1\]

OpenStudy (anonymous):

Then again, maybe not.

OpenStudy (anonymous):

I'm work at now, gonna retry your solution when I get home.

OpenStudy (amistre64):

ok, ill keep at it, since im sure i could use the practice :)

OpenStudy (amistre64):

the wolf gives an Ei function which tells me that this doesnt look like it has a pretty solution :) http://www.wolframalpha.com/input/?i=xy%22+-y%27%2Bxy%27+%2B+y+%3D+x%5E2e%5E%282x%29

OpenStudy (amistre64):

for a review if you need it, or just for nostalgia; herb gross does an awesome job at explaning power series solutions http://archive.org/details/CalculusRevisitedPart3-PowerSeriesSolutions

OpenStudy (amistre64):

he also does the other methods so the website is a good resource overall